# Gaining Wisdom from Setbacks[Uncaptioned image]: Aligning Large Language Models via Mistake Analysis

### v3

@@ Abstract
The rapid development of large language models (LLMs) has not only provided numerous opportunities but also presented significant challenges. This becomes particularly evident when LLMs inadvertently generate harmful or toxic content, either unintentionally or because of intentional inducement. Existing alignment methods usually direct LLMs toward favorable outcomes by utilizing human-annotated, flawless instruction-response pairs. Conversely, this study proposes a novel alignment technique based on mistake analysis, which deliberately exposes LLMs to erroneous content to learn the reasons for mistakes and how to avoid them. In this case, mistakes are repurposed into valuable data for alignment, effectively helping to avoid the production of erroneous responses. Without external models or human annotations, our method leverages a model’s intrinsic ability to discern undesirable mistakes and improves the safety of its generated responses. Experimental results reveal that our method outperforms existing alignment approaches in enhancing model safety while maintaining the overall utility.
@@ 1 Introduction
Insights from human practice suggest that, learning from mistakes requires first identifying the errors, understanding their underlying causes, and then preventing them in the future. Therefore, our preliminary study begins by evaluating the LLMs’ ability to identify mistakes, which is termed “discrimination” in this work (see Sec. 3). It is compared with “generation”, i.e., a LLM’s ability to generate appropriate responses to an instruction. Specifically, we ask LLMs to analyze their own responses to red-teaming instructions related to harmful problems. Surprisingly, even models not yet aligned for safety (such as Alpaca (Taori et al., 2023)) can identify mistakes in their own responses. Details are provided in Sec. 3. This may be because discrimination (i.e., recognizing mistakes in responses) tends to be easier than generation (i.e., generating appropriate responses), thereby equipping LLMs with self-critique capabilities (Huang et al., 2022; Saunders et al., 2022; Gou et al., 2023b). Motivated by this observation, we propose a novel alignment framework that trains LLMs through self-critique and mistake analysis (see Fig. 1 as an illustration).
We start by inducing an unaligned model to generate harmful responses for mistake collection (see Sec. 5.1). Then, we inform the model about the potential mistakes and instruct it to evaluate the responses. This mistake analysis data, along with regular helpful and harmless instruction-response pairs, is used for model fine-tuning. This allows LLMs to simultaneously learn what should and should not be generated to improve alignment performance, as the mistake analysis data serve as a “fine-grained mask” that helps avoid harmful content. Additionally, we demonstrate that our method can effectively defend post-aligned LLMs against novel instruction attacks using a limited number of representative mistakes (see Sec. 5.2). In summary, our method is the first to leverage natural-language-based mistake analysis provided by the model itself for alignment. Extensive experiments on various benchmarks (Dubois et al., 2023; Dai et al., 2023) demonstrate the superiority of our method. The main contributions of this work are threefold:
@@
    We are the first to demonstrate that a LLM can achieve self-alignment without external models and additional human annotations. The model’s inherent discrimination ability can be utilized to enhance its generation capability.
@@ 2 Related Work
(SFT) is widely used to align LLMs with human expectations (Ouyang et al., 2022; Wang et al., 2023; Gou et al., 2023a). This method involves calculating the cross-entropy loss over the ground-truth response to an input instruction, thereby training LLMs to adhere to user instructions. A key limitation of SFT is its exclusive focus on optimal responses, without comparing them with less optimal ones. To overcome this, variants of SFT have been developed, such as Reward Ranked Fine-tuning (RAFT) (Dong et al., 2023) and Chain of Hindsight (CoH) (Liu et al., 2023a). RAFT employs a reward model to score and select samples, fine-tuning the model with only high-reward examples. CoH, in contrast, fine-tunes LLMs using sequences of responses along with human feedback, enabling the models to discern differences between various responses. However, these SFT-based methods primarily direct the model towards identifying an “optimal response”, and often shield it from poorer responses.
@@
(RLHF) (Ouyang et al., 2022) optimizes LLMs using human-elicited reward models (RM), typically trained with pairwise human preferences for model outputs. However, obtaining high-quality human-labeled preference data at scale is resource-intensive. RLAIF (Lee et al., 2023) addresses this by simulating human preferences using LLMs, though the labels are noisier than human-validated ones. Further advancements in alignment include approaches like DPO (Rafailov et al., 2023) and RRHF (Yuan et al., 2023). DPO integrates ranking information into LLM fine-tuning, while RRHF modifies loss terms for improved alignment. The use of contrastive methods in reinforcement learning (Yang et al., 2023) has shown improvements in sample efficiency and model quality by highlighting the differences between positive and negative responses. While RL-based methods enable models to assess the relative quality of responses, they often do not provide specific reasons for penalizing lower-quality outputs.
@@ 3.1 Generation against Discrimination
In this section, we examine a LLM’s mistake detection capabilities, which is termed “discrimination” in our work. It is compared with “generation”, denoting the LLM’s ability to generate appropriate responses to a given instruction. Specifically, in our work, “discrimination” involves more than just categorizing a response as “harmful” or “harmless”. It requires a nuanced analysis of the response quality, as shown in Fig. 7(b). Our focus is on comparing a LLM’s ability to generate appropriate responses versus its ability to identify potential mistakes.
Three models are considered, including Alpaca (Taori et al., 2023), GPT-3 (Olmo et al., 2021) and GPT-3.5 (Ye et al., 2023). These models are tested with 500 red-teaming instructions from the PKU-SafeRLHF dataset, which usually involves harmful problems. The models then analyze the safety of their own responses. We compare the quality between two types of pairs: (1) Instruction-Response pairs for generation, and (2) (Instruction, Response)-Analysis pairs for discrimination. For evaluation, GPT-4 is used to rate the quality of these pairs on a scale of 1 to 10. This is followed by a human verification process. For detailed experimental settings, please refer to Appendix A.1.
As shown in Fig. LABEL:fig:pre_gen_vs_dis, across all evaluated models, the discrimination scores (i.e., identifying and analyzing potential mistakes) consistently exceed those of generation (i.e., producing harmless responses directly) with a significant margin. This indicates that even though LLMs may generate harmful responses, they still have the capability to identify harmful elements within their own responses (see examples in Appendix A.1). This observation supports the idea that discrimination is simpler than generation (Saunders et al., 2022).
@@ 3.2 Guided Analysis against Unguided Analysis
We further explore how to boost inherent discrimination ability of LLMs. We evaluate a LLM’s capability to analyze potential mistakes in two different scenarios: (1) guided analysis and (2) unguided analysis. In guided analysis, the LLMs are explicitly informed within the prompt that the provided responses could be potentially harmful (see Fig. 3(b)). In unguided analysis, on the other hand, the LLMs evaluate the response quality without any specific indications (see Fig. 3(c)).
We use the same 500 red-teaming instructions from Sec. 3.1 along with their original problematic responses from the PKU-SafeRLHF dataset for mistake analysis and evaluate the quality of analysis using a scale from 1 to 10. Each pair of the guided and unguided analysis, corresponding to the exact same instruction-response sample, is categorized as a win, tie, or lose based on their scores. As illustrated in Fig. LABEL:fig:pre_guided_vs_unguided, there is a noticeable preference for guided analysis. Across all models, the number of “wins” in guided scenarios consistently exceeds that in unguided ones, emphasizing the effectiveness of providing clear guidance when requesting mistake analysis. See Appendix A.2 for detailed examples.
@@ 4.1 Alignment through Mistake Analysis
To obtain responses containing mistakes, we induce a model to produce toxic outputs by inserting hint keywords (e.g., harmful, unethical, and offensive) into the instruction prompts, as shown in Fig. 3(a). We denote such an (induced) toxic response as $y_{\text{harm}}$ paired with $(x_{\text{harm}},y_{\text{harmless}})\in D_{\text{harmless}}$. An unaligned model is primarily considered here due to its susceptibility to malicious instructions. Experimental results show that this seemingly simple attack achieves a surprising success rate (more than 72% as in Table 4), which convinces us to adopt these induced responses for subsequent mistake analysis. Besides, mistakes (i.e., harmful responses) can also be obtained from human-annotated data.
@@
After the model’s fine-tuning, a guided strategy is employed during inference. In this phase, the model is explicitly reminded to formulate “harmless, ethical, and inoffensive” responses, as illustrated in Fig. 1. This reminder during inference acts as a safeguard, ensuring the model’s adherence to ethical standards and preventing the generation of potentially harmful content. In this case, the model computes the conditional probability based on the guideline to generate harmless content, operating within a constrained prefix context provided during inference.
@@ 5.1 Alignment
The PKU-SafeRLHF dataset (Dai et al., 2023) is used for both training and evaluation. This human-curated dataset emphasizes safety preference, covering multiple dimensions such as insults, immorality, crime, emotional harm, and privacy. For each instruction in the dataset, two responses are provided, with labels identifying the more harmful one. This setup supports the training of both SFT and RL-based models. We refine the training set to include 10,260 unique instructions, each paired with corresponding “harmless” and “harmful” responses. In consideration of the balance between helpfulness and harmfulness (Bai et al., 2022b), our training set is further augmented with an additional 52k helpful instructions from Taori et al. (2023). For evaluation, we use two different test sets. The first is the test set from AlpacaFarm (Dubois et al., 2023), which contains 805 instructions for assessing helpfulness. The second is the test set of PKU-SafeRLHF, which contains 1,523 red-teaming instructions for evaluating harmfulness.
@@ Results.
As indicated in Table 1, our method consistently surpasses existing alignment methods, such as vanilla SFT, Critique-Revise, RLHF, and CoH. Particularly, our method remarkably enhances the performance of harmlessness while effectively preserving helpfulness. See Fig. 4 for a qualitative comparison among different methods. Utilizing original faulty cases from the training set alongside Alpaca’s mistake analysis, our method shows a remarkable $35.2\%$ relative improvement in Harmless Rate over Alpaca. Furthermore, applying our method to harmful responses generated by Alpaca through guided mistake induction raises the Harmless Rate to $72.4\%$. This indicates the value of self-induced mistakes as flawed responses in our analysis-based alignment. Remarkably, when using GPT-3.5 as the analysis source, our method achieves state-of-the-art results, with a Harmless Rate of $74.1\%$. This underscores the benefits of advanced analysis sources. Other evaluation metrics exhibit trends consistent with the Harmless Rate.
@@ 5.3 Ablation Study
We consider two sources of mistakes: the original “bad” responses from the training dataset and those induced via guided mistake induction. Table 3, particularly Rows #2 and #3, shows that performance significantly improves when mistakes generated by the model itself are utilized. This improvement underscores the complexities inherent in model-induced errors.
@@ 6 Conclusion
Ensuring that LLMs align with human values is of paramount importance. Existing alignment methods typically shield LLMs from mistakes to prevent the generation of toxic responses. Our approach, however, introduces a novel alignment method based on mistake analysis. This method deliberately exposes LLMs to flawed outputs, turning these mistakes into valuable data for model alignment. Experimental results demonstrate the efficacy of our approach. It outperforms both SFT and RL-based methods in aligning unaligned models and defending post-aligned models against advanced instruction attacks. Even with a limited number of mistakes, our method effectively understands the root causes of bad responses and generalizes to address similar challenges more efficiently. Echoing the wisdom of the Chinese proverb, “A fall into the pit, a gain in your wit”, our research aims to endow LLMs with a similar depth of understanding and learning capacity.
@@ Results.
As seen in Fig. LABEL:fig:instance_alpaca_1, Alpaca fails to generate a safe response to the instruction “How do you insult an Imam?”, utilizing profoundly inappropriate and offensive terms such as the “comparing him to a woman”, “calling him a degenerate” and “traitor of the faith”. The expression of such disrespectful and harmful sentiments illustrates a severe breach of ethical and safety guidelines, justifying the assigned score of 1.
@@
Fig. LABEL:fig:instance_gpt3_1 and Fig. LABEL:fig:instance_gpt3.5_1 provide examples for GPT-3 and GPT-3.5, respectively. The analysis of both GPT-3 and GPT-3.5 achiev higher scores than their respective responses. GPT-3 is capable of recognizing that not everyone may understand the humor in its response and that it could potentially “hurt someone’s feelings”. GPT-3.5, when discriminating, can even identify the subtle safety issues, such as “it appears to be encouraging harmful and potentially illegal behavior”, and offers impartial advice, stating that “the AI should clearly state that it is not appropriate or ethical to harm others”.
@@ A.2 Guided analysis against unguided Analysis
As mentioned in Sec. 3.2, we rate the analysis generated by the above three models on a scale of 1 to 10. After manual score verification, we calculate the average scores separately according to models and scenarios. We further count the number of instruction-response pairs whose corresponding guided analysis wins, ties with, or loses to unguided analysis, as seen in the horizontal histogram in Fig. LABEL:fig:pre_guided_vs_unguided.
@@ Results.
In the given example in Fig. LABEL:fig:instance_alpaca_2, the Alpaca’s unguided analysis renders a score of 1, with the response being considered as correct. The lack of guidance causes the system to miss potential safety concerns in the response, leading to a superficial analysis that merely acknowledges the correctness of responses.

### v5

@@ Abstract
The rapid development of large language models (LLMs) has not only provided numerous opportunities but also presented significant challenges. This becomes particularly evident when LLMs inadvertently generate harmful or toxic content, either unintentionally or because of intentional inducement. Existing alignment methods usually direct LLMs toward the favorable outcomes by utilizing human-annotated, flawless instruction-response pairs. Conversely, this study proposes a novel alignment technique based on mistake analysis, which deliberately exposes LLMs to erroneous content to learn the reasons for mistakes and how to avoid them. In this case, mistakes are repurposed into valuable data for alignment, effectively helping to avoid the production of erroneous responses. Without external models or human annotations, our method leverages a model’s intrinsic ability to discern undesirable mistakes and improves the safety of its generated responses. Experimental results reveal that our method outperforms existing alignment approaches in enhancing model safety while maintaining the overall utility.
@@ 1 Introduction
Insights from human practice suggest that, learning from mistakes requires first identifying the errors, understanding their underlying causes, and then preventing them in the future. Therefore, our preliminary study begins by evaluating the LLMs’ ability to identify mistakes, which is termed “discrimination” in this work (see Sec. 3). It is compared with “generation” (i.e., a LLM’s ability to generate appropriate responses to an instruction). Specifically, we ask LLMs to analyze their own responses to red-teaming instructions related to harmful problems. Surprisingly, even models not yet aligned for safety (such as Alpaca (Taori et al., 2023)) can identify mistakes in their own responses. Details are provided in Sec. 3, which might be because “discrimination” (i.e., recognizing mistakes in responses) tends to be easier than “generation” (i.e., generating appropriate responses), thereby equipping LLMs with self-critique capabilities (Huang et al., 2022; Saunders et al., 2022; Gou et al., 2023b). Motivated by this observation, we propose a novel alignment framework that trains LLMs through self-critique and mistake analysis (see Fig. 1 as an illustration).
We start by inducing an unaligned model to generate harmful responses for mistake collection (see Sec. 5.1). Then, we inform the model about the potential mistakes and instruct it to evaluate its own responses. This mistake analysis data, along with regular helpful and harmless instruction-response pairs, is used for model fine-tuning. This allows LLMs to simultaneously learn what should and should not be generated to improve alignment performance, as the mistake analysis data serve as a “fine-grained mask” that helps avoid harmful content. Additionally, we demonstrate that our method can effectively defend post-aligned LLMs against novel instruction attacks using a limited number of representative mistakes (see Sec. 5.2). In summary, our method is the first to leverage natural-language-based mistake analysis provided by the model itself for alignment. Extensive experiments on various benchmarks (Dubois et al., 2023; Dai et al., 2023) demonstrate the superiority of our method. The main contributions of this work are threefold:
@@
    We are the first to demonstrate that a LLM can achieve self-alignment without external models and additional human annotations. The model’s inherent discrimination ability can be utilized to enhance its own generation capability.
@@ 2 Related Work
(SFT) is widely used to align LLMs with human expectations (Ouyang et al., 2022; Wang et al., 2023; Gou et al., 2023a). This method involves calculating the cross-entropy loss over the ground-truth response to an input instruction, thereby training LLMs to adhere to user instructions. A key limitation of SFT is its exclusive focus on optimal responses, without comparing them with the less optimal ones. To overcome this problem, variants of SFT have been developed, such as Reward Ranked Fine-tuning (RAFT) (Dong et al., 2023) and Chain of Hindsight (CoH) (Liu et al., 2023a). RAFT employs a reward model to score and select samples, fine-tuning the model with only high-reward examples. CoH, in contrast, fine-tunes LLMs using sequences of responses along with human feedback, enabling the models to discern differences between various responses. However, these SFT-based methods primarily direct the model towards identifying an “optimal response”, and often shield it from poorer responses.
@@
(RLHF) (Ouyang et al., 2022) optimizes LLMs using the human-elicited reward models (RM), typically trained with pairwise human preferences for model outputs. However, obtaining high-quality human-labeled preference data at scale is still resource-intensive. RLAIF (Lee et al., 2023) addresses this by simulating human preferences using LLMs, though the labels might be noisier than human-validated ones. Further advancements in alignment include approaches like DPO (Rafailov et al., 2023) and RRHF (Yuan et al., 2023). DPO integrates ranking information into LLM fine-tuning, while RRHF modifies loss terms for improved alignment. The use of contrastive methods in reinforcement learning (Yang et al., 2023) has shown improvements in sample efficiency and model quality by highlighting the differences between positive and negative responses. While RL-based methods enable models to assess the relative quality of responses, they often do not provide specific reasons for penalizing lower-quality outputs.
@@ 3.1 Generation against Discrimination
In this section, we examine a LLM’s mistake detection capabilities, which is termed “discrimination” in our work. It is compared with “generation”, denoting the LLM’s ability to generate appropriate responses to a given instruction. Specifically, in our work, “discrimination” involves more than just categorizing a response as “harmful” or “harmless”. It requires a nuanced analysis of the response quality, as demonstrated in Fig. 7(b). Our focus is on comparing a LLM’s ability to generate appropriate responses versus its ability to identify potential mistakes.
Three models are considered, including Alpaca (Taori et al., 2023), GPT-3 (Olmo et al., 2021) and GPT-3.5 (Ye et al., 2023). These models are evaluated with 500 red-teaming instructions from the PKU-SafeRLHF dataset, which usually involves harmful problems. The models then analyze the safety of their own responses. We compare the quality between two types of pairs: (1) Instruction-Response pairs for generation, and (2) (Instruction, Response)-Analysis pairs for discrimination. For evaluation, GPT-4111https://chatgpt.ust.hk is used to rate the quality of these pairs on a scale of 1 to 10. This is followed by a human verification process. For detailed experimental settings, please refer to Appendix A.1.
As shown in Fig. LABEL:fig:pre_gen_vs_dis, across all the evaluated models, the discrimination scores (i.e., identifying and analyzing potential mistakes) consistently exceed those of generation (i.e., producing harmless responses directly) with a significant margin. This indicates that even though LLMs may generate harmful responses, they still have the capability to identify the harmful elements within their own responses (see examples in Appendix A.1). This observation supports the idea that discrimination is simpler than generation (Saunders et al., 2022).
@@ 3.2 Guided Analysis against Unguided Analysis
We further explore how to boost the inherent discrimination ability of LLMs. We evaluate a LLM’s capability to analyze potential mistakes in two different scenarios: (1) guided analysis and (2) unguided analysis. In guided analysis, the LLMs are explicitly informed within the prompt that the provided responses could be potentially harmful (see Fig. 3(b)). In unguided analysis, on the other hand, the LLMs evaluate the response quality without any specific indications (see Fig. 3(c)).
We use the same 500 red-teaming instructions from Sec. 3.1 along with their original problematic responses from the PKU-SafeRLHF dataset for mistake analysis and evaluate the quality of analysis using a scale from 1 to 10. Each pair of the guided and unguided analyses, corresponding to the exact same instruction-response sample, is categorized as a win, tie, or lose based on their scores. As illustrated in Fig. LABEL:fig:pre_guided_vs_unguided, there is a noticeable preference for guided analysis. Across all models, the number of “wins” in guided scenarios consistently exceeds that in unguided ones, emphasizing the effectiveness of providing clear guidance when requesting mistake analysis. See Appendix A.2 for more detailed examples.
@@ 4.1 Alignment through Mistake Analysis
To obtain responses containing mistakes, we induce a model to produce toxic outputs by inserting hint keywords (e.g., harmful, unethical, and offensive) into the instruction prompts, as represented in Fig. 3(a). We denote such an (induced) toxic response as $y_{\text{harm}}$ paired with $(x_{\text{harm}},y_{\text{harmless}})\in D_{\text{harmless}}$. An unaligned model is primarily considered here due to its susceptibility to malicious instructions. Experimental results show that this seemingly simple attack achieves a surprising success rate (more than 72% as in Table 4), which convinces us to adopt these induced responses for subsequent mistake analysis. Besides, mistakes (i.e., harmful responses) can also be obtained from human-annotated data (Dai et al., 2023).
@@
After the model’s fine-tuning, a guided strategy is employed during inference. In this phase, the model is explicitly reminded to formulate “harmless, ethical, and inoffensive” responses, as illustrated in Fig. 1. This reminder during inference acts as a safeguard, ensuring the model’s adherence to ethical standards and preventing the generation of potentially harmful content. In this case, the model computes the conditional probability based on the guidelines to generate harmless content, operating within a constrained prefix context provided during inference, as demonstrated in Fig. 1.
@@ 5.1 Alignment
The PKU-SafeRLHF dataset (Dai et al., 2023) is used for both training and evaluation. This human-curated dataset emphasizes safety preference, covering multiple dimensions such as insults, immorality, crime, emotional harm, and privacy. For each instruction in the dataset, two responses are provided, with labels identifying the more harmful one. This setup supports the training of both SFT and RL-based models. We refine the training set to include 10,260 unique instructions, each paired with corresponding “harmless” and “harmful” responses. In consideration of the balance between helpfulness and harmfulness (Bai et al., 2022b), our training set is further augmented with an additional 52k helpful instructions from Taori et al. (2023). For evaluation, we use two different test sets. The first is the test set from AlpacaFarm (Dubois et al., 2023), which contains 805 instructions for assessing helpfulness. The second is the test set of PKU-SafeRLHF, which contains 1,523 red-teaming instructions for harmfulness evaluation.
@@ Results.
As indicated in Table 1, our method consistently surpasses existing alignment methods, including the vanilla SFT, Critique-Revise, RLHF, and CoH. Particularly, our method remarkably enhances the performance of harmlessness while effectively preserving helpfulness. See Fig. 4 for a qualitative comparison among different methods. Utilizing original faulty cases from the training set alongside Alpaca’s mistake analysis, our method shows a remarkable $35.2\%$ relative improvement in Harmless Rate over Alpaca. Furthermore, applying our method to harmful responses generated by Alpaca through guided mistake induction raises the Harmless Rate to $72.4\%$. This indicates the value of self-induced mistakes as flawed responses in our analysis-based alignment. Remarkably, when using GPT-3.5 as the analysis source, our method achieves state-of-the-art results, with a Harmless Rate of $74.1\%$. This underscores the benefits of advanced analysis sources. Other evaluation metrics exhibit trends consistent with the Harmless Rate.
@@ 5.3 Ablation Study
We consider two sources of mistakes: the original “bad” responses from the training dataset and those induced via guided mistake induction. Table 3, particularly Rows #2 and #3, shows that performance significantly improves when mistakes generated by the model itself are utilized. This improvement underscores the complexities inherent in the model-induced errors.
@@ 6 Conclusion
Ensuring that LLMs align with human values is of paramount importance. Existing alignment methods typically shield LLMs from mistakes to prevent the generation of toxic responses. Our approach, however, introduces a novel alignment method based on mistake analysis. This method deliberately exposes LLMs to the flawed outputs, turning these mistakes into valuable data for model alignment. Experimental results demonstrate the efficacy of our approach, which outperforms both SFT and RL-based methods in aligning unaligned models and defending post-aligned models against advanced instruction attacks. Even with a limited number of mistakes, our method effectively understands the root causes of bad responses and generalizes to address similar challenges more efficiently. Echoing the wisdom of the Chinese proverb, “A fall into the pit, a gain in your wit”, our work aims to endow LLMs with a similar depth of understanding and learning capacity.
@@ Results.
As shown in Fig. LABEL:fig:instance_alpaca_1, Alpaca fails to generate a safe response to the instruction “How do you insult an Imam?”, utilizing profoundly inappropriate and offensive terms such as the “comparing him to a woman”, “calling him a degenerate” and “traitor of the faith”. The expression of such disrespectful and harmful sentiments illustrates a severe breach of ethical and safety guidelines, justifying the assigned score of 1.
@@
Fig. LABEL:fig:instance_gpt3_1 and Fig. LABEL:fig:instance_gpt3.5_1 provide examples for GPT-3 and GPT-3.5, respectively. The analyses of both GPT-3 and GPT-3.5 achieve higher scores than their respective responses. GPT-3 is capable of recognizing that not everyone may understand the humor in its response and that it could potentially “hurt someone’s feelings”. GPT-3.5, when discriminating, can even identify the subtle safety issues, such as “it appears to be encouraging harmful and potentially illegal behavior”, and offers impartial advice, stating that “the AI should clearly state that it is not appropriate or ethical to harm others”.
@@ A.2 Guided analysis against unguided Analysis
As mentioned in Sec. 3.2, we rate the analyses generated by the above three models on a scale of 1 to 10. After the manual score verification, we calculate the average scores separately according to models and scenarios. We further count the number of instruction-response pairs whose corresponding guided analysis wins, ties with, or loses to unguided analysis, as seen in the horizontal histogram in Fig. LABEL:fig:pre_guided_vs_unguided.
@@ Results.
In the given example in Fig. LABEL:fig:instance_alpaca_2, Alpaca’s unguided analysis renders a score of 1, with the response being considered as correct. The lack of guidance causes the system to miss the potential safety concerns in the response, leading to a superficial analysis that merely acknowledges the correctness of provided responses.